# Momentum-SAM: Sharpness Aware Minimization without Computational Overhead

### v2

@@ 1.1 Related Work
Unlike the aforementioned methods, several algorithms were proposed to reduce SAMs runtime, mostly sharing the idea of reducing the number of additional forward/backward passes, in contrast to our approach which relies on finding more efficient parameter perturbations. For example, Jiang et al. (2023) are evaluating in each iteration if a perturbation calculation is to be performed. LookSAM (Liu et al., 2022) updates perturbations only each $k$-th iterations and applies perturbation components orthogonal to SGD gradients in iterations in between. Mi et al. (2022) are following an approach based on sparse matrix operations and ESAM (Du et al., 2022b) combines parameter sparsification with the idea to reduce the number of input samples for second forward/backward passes. Similarly, Bahri et al. (2021) and Ni et al. (2022) calculate perturbations on micro-batches. Not explicitly targeted at efficiency optimization, Mueller et al. (2023) show that only perturbing Batch Norm layers even further improves SAM. SAF and its memory efficient version MESA were proposed by Du et al. (2022a), focusing on storing past iterations weights to minimize sharpness on the digits output instead of the loss function. Perturbations in momentum direction after the momentum buffer update resulting in better performance but no speedup where proposed by Li and Giannakis (2023). Furthermore, several concepts were proposed to explain the success of SAM and related approaches going beyond sharpness reduction of the training loss landscape (Andriushchenko and Flammarion, 2022; Möllenhoff and Khan, 2023; Andriushchenko et al., 2023b), with the influence on the balance of features/activations being a promising alternative explanation (Andriushchenko et al., 2023a; Springer et al., 2024), which we also investigate in Appx. A.5.
@@ 2.1 Notation
Given a finite training dataset $\mathcal{S}\subset\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}$ is the set of possible inputs and $\mathcal{Y}$ the set of possible targets drawn from a joint distribution $\mathfrak{D}$, we study a model $f_{\bm{w}}:\mathcal{X}\rightarrow\mathcal{Y}$ parameterized by $\bm{w}\in\mathcal{W}$, an element-wise loss function $l:\mathcal{W}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the distribution loss $L_{\mathfrak{D}}(\bm{w})=\mathbb{E}_{(x,y)\sim\mathfrak{D}}(l(\bm{w},x,y))$ and the empirical (training) loss $L_{\mathcal{S}}(\bm{w})=1/|\mathcal{S}|\sum_{(x,y)\in\mathcal{S}}l(\bm{w},x,y)$. If calculated on a single batch $\mathcal{B}\subset S$ we denote the loss as $L_{\mathcal{B}}$. We denote the L2-norm by |||||$\cdot$||| | ⋅ | |.
@@ 2.2 Sharpness Aware Minimization (SAM)
$\underset{\bm{w}}{\text{min}}~{}\underset{||\bm{\epsilon}||\leq\rho}{\text{max}}~{}L_{\mathcal{S}}(\bm{w}+\bm{\epsilon})$ (1)
@@
$L_{\mathcal{B}}^{\text{SAM}}(\bm{w})\coloneqq L_{\mathcal{B}}(\bm{w}+\bm{\epsilon}^{\text{SAM}})~{}~{}\text{where}~{}~{}\bm{\epsilon}^{\text{SAM}}\coloneqq\rho\frac{\nabla L_{\mathcal{B}}(\bm{w})}{||\nabla L_{\mathcal{B}}(\bm{w})||}.$ (2)
@@ 2.4 Momentum-SAM
Foret et al. (2021) show that performing multiple iterations of projected gradient ascent in the inner maximization does result in parameters with higher loss inside the $\rho$-ball (cf. Eq. 1). However, and counterintuitively, this improved inner maximization does not yield a better generalization of the model. We conclude that finding the exact (per batch) local maximum is not pivotal to SAM. Inspired by NAG and given that the theoretical framework of sharpness minimization is based on calculating the sharpness on the full training dataset, we propose using the momentum vector as the perturbation direction and call the resulting algorithm Momentum-SAM (MSAM) (further perturbations are discussed in Appx. A.6). Following the above notation, this yields the loss objective
$L_{\mathcal{B}}^{\text{MSAM}}(\bm{w})\coloneqq L_{\mathcal{B}}(\bm{w}+\bm{\epsilon}^{\text{MSAM}})~{}~{}\text{where}~{}~{}\bm{\epsilon}^{\text{MSAM}}\coloneqq-\rho\frac{\bm{v}_{t}}{||\bm{v}_{t}||}.$ (5)
@@
For an efficient implementation, we shift the starting point of each iteration to be the perturbed parameters $\bm{\widetilde{w}}_{t}=\bm{w}_{t}-\rho\bm{v}_{t}/||\bm{v}_{t}||$ (in analogy to common implementations of NAG) and remove the final perturbation after the last iteration (see Alg. 1). All mentioned optimization strategies are depicted in detail in Fig. 1. Since SGD with momentum as well as Adam store a running mean of gradients, MSAM does not take up additional memory and comes with negligible computational overhead.
@@
      $\bm{w}_{t}=\bm{\widetilde{w}}_{t}+{\color[rgb]{0,0.390625,0}\rho\frac{\bm{v}_{t}}{||\bm{v}_{t}||}}$ // remove last pert.
      $\bm{v}_{t+1}$ $=$ $\mu\bm{v}_{t}$ $+$ $\bm{g}_{\text{MSAM}}$ // update momentum
@@ 3.1 Speed and Accuracy for ResNets on CIFAR100
  SAM $84.16$±$\pm 0.12$ $79.25$±$\pm 0.10$ $83.36$±$\pm 0.17$ $69.1$ 0.52
  Baseline $81.51$±$\pm 0.09$ $76.90$±$\pm 0.15$ $81.46$±$\pm 0.13$ $67.0$ $\bm{1.00}$
NAG $82.00$±$\pm 0.11$ $77.09$±$\pm 0.18$ $82.12$±$\pm 0.12$- $\bm{0.99}$
LookSAM $\bm{83.31}$±$\pm 0.12$ $\bm{79.00}$±$\pm 0.08$ $82.24$±$\pm 0.11$ $68.0$ 0.84
ESAM $82.71$±$\pm 0.38$ $77.79$±$\pm 0.11$ $80.49$±$\pm 0.40$ $66.1$ 0.62
MESA $82.75$±$\pm 0.08$ $78.32$±$\pm 0.08$ $81.94$±$\pm 0.26$ $\bm{69.0}$ 0.77
MSAM (ours) $\bm{83.21}$±$\pm 0.07$ $\bm{79.11}$±$\pm 0.09$ $\bm{82.65}$±$\pm 0.12$ $\bm{69.1}$ $\bm{0.99}$
@@ 3.2 ResNet and ViT on ImageNet Results
In addition, we conducted experiments for ViT-S/32 on ImageNet when giving MSAM the same computational budget as SAM (i.e. training for 180 epochs) yielding a test accuracy of 70.1% and thus clearly outperforming SAMs 69.1% (also see Appx. A.8).
Table 2: Test accuracies on ImageNet for baseline optimizers (SGD or AdamW), SAM and MSAM. Estimated uncertainties: ResNet: ±$\pm 0.08$, ViT (90 epochs): ±$\pm 0.17$, ViT (300 epochs): ±$\pm 0.24$. Improvements over baseline are given in green. MSAM yields results comparable to SAM for most models while being $\approx 2$ times faster in all our experiments.
@@
  ResNet-50 100 SGD $76.3$ $76.6$$+0.3$ $76.5$$+0.2$
ResNet-101 100 SGD $77.9$ $78.7$$+0.8$ $78.2$$+0.3$
ViT-S/32 300 AdamW $67.2$ $71.4$$+4.2$ $70.5$+3.33.3+3.3+ 3.3
90 AdamW $67.0$ $69.1$$+2.1$ $69.1$$+2.1$
ViT-S/16 300 AdamW $73.0$ $78.2$$+5.2$ $75.8$$+2.8$
90 AdamW $72.6$ $75.8$$+3.2$ $74.9$$+2.3$
ViT-B/32 90 AdamW $66.9$ $70.4$$+3.5$ $70.2$+3.33.3+3.3+ 3.3
ViT-B/16 90 AdamW $73.0$ $77.7$$+4.7$ $75.7$$+2.7$
@@ 4 Properties of MSAM
In the following we analyze the perturbation direction of MSAM, its similarities to SAM and the loss sharpness in more detail. Additional experiments are presented in the appendix. These include the observations that the perturbation normalization is not pivotal to MSAM and instead the binding to the learning rate $\eta$ discriminates NAG and MSAM (Appx. A.2) and that MSAM reduces feature ranks more effectively than SAM (Appx. A.5).
@@
                                         SGD $81.51$±$\pm 0.09$ $76.90$±$\pm 0.15$ $81.46$±$\pm 0.13$
vanilla SAM $84.16$±$\pm 0.12$ $79.25$±$\pm 0.10$ $83.36$±$\pm 0.17$
MSAM $83.21$±$\pm 0.07$ $79.11$±$\pm 0.09$ $82.65$±$\pm 0.12$
adaptive SAM $84.74$±$\pm 0.13$ $79.96$±$\pm 0.13$ $83.30$±$\pm 0.06$
MSAM $84.15$±$\pm 0.13$ $79.89$±$\pm 0.09$ $83.48$±$\pm 0.08$
BN-only SAM $84.57$±$\pm 0.07$ $79.73$±$\pm 0.24$ $84.51$±$\pm 0.17$
MSAM $83.62$±$\pm 0.09$ $79.73$±$\pm 0.14$ $83.49$±$\pm 0.19$
@@ 4.2 Similarity between SAM and MSAM
$\bm{g}_{\text{SAM}}$ $=\nabla L_{\mathcal{B}_{t}}(\bm{w}_{t}+\rho_{\text{SAM}}\nabla L_{\mathcal{B}_{t}}(\bm{w}_{t})/||\nabla L_{\mathcal{B}_{t}}(\bm{w}_{t})||)$
$\bm{g}_{\text{MSAM}}(\rho_{\text{MSAM}})$ $=\nabla L_{\mathcal{B}_{t}}(\bm{w}_{t}-\rho_{\text{MSAM}}\bm{v}_{t}/||\bm{v}_{t}||)$
@@ 4.3 Loss Sharpness Analysis
In Fig. 4C, we choose filter-normalized random vectors as perturbations as in Li et al. (2018). Since the loss landscape is rotational symmetric around the origin for multiple directions (as used in the original paper), we confine our analysis to one perturbation direction allowing an easier comparison (see Appx. A.11 for two perturbation directions). MSAM reaches the lowest sharpness, while both MSAM and SAM, significantly flatten the loss. This might be caused by MSAM approximating the maximization of $L_{\mathcal{S}}$ better due to the momentum vector $\bm{v}_{t}$ being an aggregation of gradients of multiple batches. Interestingly, this relates to the findings of Foret et al. (2021) regarding $m$-sharpness, where the authors performed the maximization on even smaller data samples (fractions of batches per GPU in distributed training), yielding even better generalization. In this sense MSAM, reduces $m$ even further over ordinary SAM ($m=1$). In the same line of argument and despite being more efficient, MSAM oftentimes does not improve generalization compared to SAM. However, contradicting the general idea behind correlations of generalization and sharpness, MSAM yields flatter minima (if defined as in Fig. 4C), hence, indicating that explanations for the improved generalization of SAM/MSAM go beyond the reduction in sharpness. Additionally, we show that SAM guides the optimization through regions of lowest curvature compared to SGD and SAM in Appx. A.4.
@@ A.4 Curvature
We calculated the loss curvature in momentum direction, gradient direction and in random direction in Fig. A.8 if training with SGD, SAM and MSAM for WRN-16-4 on CIFAR100. For this we calculate $\bm{\epsilon}^{T}\mathrm{Hess}(L_{\mathcal{S}}(\bm{w}))\bm{\epsilon}$ for direction vectors $\bm{\epsilon}$ (normalized to $||\bm{\epsilon}||_{2}=1$) every 50 optimizer steps.
The curvatures in momentum directions are larger than the curvature random direction (which tends towards the mean curvature as amount of parameters increase) for all optimizers and epochs, validating Setting 1 in Appx. A.1 and thus the suitability of momentum directions for sharpness estimation (especially compared to random perturbations; cf. Appx. A.6).
@@
A.5 Feature Rank Analysis
@@
A.6 Random and Last Gradient Perturbations
@@ A.7 Hyperparameter Stability
A.8 Comparison with Same Computational Budgets
@@
A.9 Machine Translation
Machine translation results (English to Romanian) on the WMT 2016 (Bojar et al., 2016)dataset by finetuning a T5-tiny model (efficient version by Tay et al. (2022)) from a publicly available checkpoint pretrained on the C4 dataset (Raffel et al., 2019). We scanned $\rho\in\{0.01,0.03,0.1,0.3,1\}$ and found MSAM and SAM both to perform best at $\rho=0.1$. The resulting BLEU scores are shown in Tab. A.1. MSAM slightly outperforms SAM (as well as AdamW) while requiring two times less computations.
Table A.1: BLEU scores for T5-tiny trained on English to Romanian translation on the WMT 2016 dataset.
AdamW SAM MSAM
$23.35$ $23.57$ $23.64$
@@ A.10 Loss Ascent in Momentum Direction
Figure A.12: Loss before $(L_{\mathcal{B}_{t}}(\bm{w}_{t}))$ and after $(L_{\mathcal{B}_{t}}(\bm{w}_{t}-\rho\bm{v}_{t}/||\bm{v}_{t}||))$ perturbation in momentum direction as done by MSAM ($\rho=3$) for WRN 16-4 on CIFAR100 for different learning rates.

### v3

@@ 1.1 Related Work
Unlike the aforementioned methods, several algorithms were proposed to reduce SAMs runtime, mostly sharing the idea of reducing the number of additional forward/backward passes, in contrast to our approach which relies on finding more efficient parameter perturbations. For example, Jiang et al. (2023) are evaluating in each iteration if a perturbation calculation is to be performed. LookSAM (Liu et al., 2022) updates perturbations only each $k$-th iterations and applies perturbation components orthogonal to SGD gradients in iterations in between. Mi et al. (2022) are following an approach based on sparse matrix operations and ESAM (Du et al., 2022b) combines parameter sparsification with the idea to reduce the number of input samples for second forward/backward passes. Similarly, Bahri et al. (2021) and Ni et al. (2022) calculate perturbations on micro-batches. Not explicitly targeted at efficiency optimization, Mueller et al. (2023) show that only perturbing Batch Norm layers even further improves SAM. SAF and its memory efficient version MESA were proposed by Du et al. (2022a), focusing on storing past iterations weights to minimize sharpness on the digits output instead of the loss function. Perturbations in momentum direction after the momentum buffer update resulting in better performance but no speedup where proposed by Li and Giannakis (2023). Furthermore, several concepts were proposed to explain the success of SAM and related approaches going beyond sharpness reduction of the training loss landscape (Andriushchenko and Flammarion, 2022; Möllenhoff and Khan, 2023; Andriushchenko et al., 2023b), with the influence on the balance of features/activations being a promising alternative explanation (Andriushchenko et al., 2023a; Springer et al., 2024), which we also investigate in Appx. A.6. SAM_and_robust analyze how SAM influences adversarial robustness, a question we explore for MSAM in Appendix A.16.
@@ 2.1 Notation
Given a finite training dataset $\mathcal{S}\subset\mathcal{X}\times\mathcal{Y}$ where $\mathcal{X}$ is the set of possible inputs and $\mathcal{Y}$ the set of possible targets drawn from a joint distribution $\mathfrak{D}$, we study a model $f_{\bm{w}}:\mathcal{X}\rightarrow\mathcal{Y}$ parameterized by $\bm{w}\in\mathcal{W}$, an element-wise loss function $l:\mathcal{W}\times\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$, the distribution loss $L_{\mathfrak{D}}(\bm{w})=\mathbb{E}_{(x,y)\sim\mathfrak{D}}(l(\bm{w},x,y))$ and the empirical (training) loss $L_{\mathcal{S}}(\bm{w})=1/|\mathcal{S}|\sum_{(x,y)\in\mathcal{S}}l(\bm{w},x,y)$. If calculated on a single batch $\mathcal{B}\subset S$ we denote the loss as $L_{\mathcal{B}}$. We denote the L2-norm by $||\cdot||$.
@@ 2.2 Sharpness Aware Minimization (SAM)
[…]
@@
[…]
@@ 2.4 Momentum-SAM
Foret et al. (2021) show that performing multiple iterations of projected gradient ascent in the inner maximization does result in parameters with higher loss inside the $\rho$-ball (cf. Eq. 1). However, and counterintuitively, this improved inner maximization does not yield a better generalization of the model. We conclude that finding the exact (per batch) local maximum is not pivotal to SAM. Inspired by NAG and given that the theoretical framework of sharpness minimization is based on calculating the sharpness on the full training dataset, we propose using the momentum vector as the perturbation direction and call the resulting algorithm Momentum-SAM (MSAM) (further perturbations are discussed in Appx. A.7). Following the above notation, this yields the loss objective
[…]
@@
[…]
@@
[…]
   $\bm{v}_{t+1}$ == $\mu\bm{v}_{t}$ ++ $\bm{g}_{\text{MSAM}}$ // update momentum
@@ 3.1 Speed and Accuracy for ResNets on CIFAR100
[…]
MSAM (ours) $\bm{83.21}$$\pm 0.07$ $\bm{79.11}$$\pm 0.09$ $\bm{82.65}$$\pm 0.12$ $\bm{69.1}$ $\bm{0.99}$
@@ 3.2 ResNet and ViT on ImageNet Results
In addition, we conducted experiments for ViT-S/32 on ImageNet when giving MSAM the same computational budget as SAM (i.e. training for 180 epochs) yielding a test accuracy of 70.1% and thus clearly outperforming SAMs 69.1% (also see Appx. A.9).
Table 2: Test accuracies on ImageNet for baseline optimizers (SGD or AdamW), SAM and MSAM. Estimated uncertainties: ResNet: $\pm 0.08$, ViT (90 epochs): $\pm 0.17$, ViT (300 epochs): $\pm 0.24$. Improvements over baseline are given in green. MSAM yields results comparable to SAM for most models while being $\approx 2$ times faster in all our experiments.
@@
  ResNet-50 100 SGD $76.3$ 76.676.6+0.3+0.3 76.576.5+0.2+0.2
ResNet-101 100 SGD $77.9$ 78.778.7+0.8+0.8 78.278.2+0.3+0.3
ViT-S/32 300 AdamW $67.2$ 71.471.4+4.2+4.2 70.570.5+3.3+3.3
90 AdamW $67.0$ 69.169.1+2.1+2.1 69.169.1+2.1+2.1
ViT-S/16 300 AdamW $73.0$ 78.278.2+5.2+5.2 75.875.8+2.8+2.8
90 AdamW $72.6$ 75.875.8+3.2+3.2 74.974.9+2.3+2.3
ViT-B/32 90 AdamW $66.9$ 70.470.4+3.5+3.5 70.270.2+3.3+3.3
ViT-B/16 90 AdamW $73.0$ 77.777.7+4.7+4.7 75.775.7+2.7+2.7
@@ 4 Properties of MSAM
In the following we analyze the perturbation direction of MSAM, its similarities to SAM and the loss sharpness in more detail. Additional experiments are presented in the appendix. These include the observations that the perturbation normalization is not pivotal to MSAM and instead the binding to the learning rate $\eta$ discriminates NAG and MSAM (Appx. A.2) and that MSAM reduces feature ranks more effectively than SAM (Appx. A.6).
@@
[…]
@@ 4.2 Similarity between SAM and MSAM
[…]
@@ 4.3 Loss Sharpness Analysis
In Fig. 4C, we choose filter-normalized random vectors as perturbations as in Li et al. (2018). Since the loss landscape is rotational symmetric around the origin for multiple directions (as used in the original paper), we confine our analysis to one perturbation direction allowing an easier comparison (see Appx. A.11 for two perturbation directions). MSAM reaches the lowest sharpness, while both MSAM and SAM, significantly flatten the loss. This might be caused by MSAM approximating the maximization of $L_{\mathcal{S}}$ better due to the momentum vector $\bm{v}_{t}$ being an aggregation of gradients of multiple batches. Interestingly, this relates to the findings of Foret et al. (2021) regarding $m$-sharpness, where the authors performed the maximization on even smaller data samples (fractions of batches per GPU in distributed training), yielding even better generalization. In this sense MSAM, reduces $m$ even further over ordinary SAM ($m=1$). In the same line of argument and despite being more efficient, MSAM oftentimes does not improve generalization compared to SAM. However, contradicting the general idea behind correlations of generalization and sharpness, MSAM yields flatter minima (if defined as in Fig. 4C), hence, indicating that explanations for the improved generalization of SAM/MSAM go beyond the reduction in sharpness. Additionally, we show that SAM guides the optimization through regions of lowest curvature compared to SGD and SAM in Appx. A.4 and yields low Hessian eigenvalues Appx. A.5.
@@ 5 Conclusion
Acknowledgements
This work was funded by the German Research Foundation (DFG CRC 1459 Intelligent Matter - Project-ID 433682494).
@@ A.4 Curvature
We calculated the loss curvature in momentum direction, gradient direction and in random direction in Fig. A.8 if training with SGD, SAM and MSAM for WRN-16-4 on CIFAR100. For this we calculate $\bm{\epsilon}^{T}\mathrm{Hess}(L_{\mathcal{S}}(\bm{w}))\bm{\epsilon}$ for direction vectors $\bm{\epsilon}$ (normalized to $\bm{\epsilon}||_{2}=1$) every 50 optimizer steps.
The curvatures in momentum directions are larger than the curvature random direction (which tends towards the mean curvature as amount of parameters increase) for all optimizers and epochs, validating Setting 1 in Appx. A.1 and thus the suitability of momentum directions for sharpness estimation (especially compared to random perturbations; cf. Appx. A.7).
@@
A.5 Eigenspectrum Analysis
We calculated the top 5 eigenvalues of the Hessian using the Lanczos algorithm for WideResNet-16-4 trained by SGD, SAM and MSAM on CIFAR-100.
Table A.1: Top 5 Hessian eigenvalues.
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$
SGD 15.19 13.79 13.23 9.98 9.23
MSAM 4.15 3.03 2.30 1.88 1.81
SAM 5.18 4.85 4.43 3.72 3.36
While SAM already effectively reduces the eigenvalues of the Hessian (and thus the curvature), MSAM results in even lower Hessian eigenvalues. These findings align with the other sharpness metrics discussed and further support the conclusion that MSAM leads to lower sharpness than SAM (and SGD).
A.6 Feature Rank Analysis
@@
A.7 Random and Last Gradient Perturbations
@@ A.8 Hyperparameter Stability
A.9 Comparison with Same Computational Budgets
@@ A.10 Loss Ascent in Momentum Direction
[…]
@@ A.15 Full Hyper Parameter Search Results
A.16 Adversarial Robustness
We evaluated the adversarial robustness of MSAM under attacks by PGD (PGD), FGSM FGSM and APGDT APGDT. We test several attack methods on a WideResNet-16-4 trained on CIFAR-100. We used default hyperparameters and $\sigma=0.1$ for gaussian noise and $\epsilon=1/255$ for the other attack methods. Results shown in A.4 were averaged over 3 randomly initialized runs per optimizer. MSAM improves adversarial robustness similarly to SAM. In particular, under an undirected Gaussian noise attack, MSAM slightly outperforms SAM, further supporting its effectiveness in sharpness minimization. These findings are consistent with those of SAM_and_robust.
Table A.4: Adversarial Robustness of MSAM compared to SAM and SGD
Gaussian Noise PGD FGSM APGDT
SGD 22.55 24.05 31.80 20.72
MSAM 26.86 28.92 35.46 27.06
SAM 25.77 30.22 35.30 26.08